# Crypto Inverse-Power Options and Fractional Stochastic Volatility

### v2

@@ 3 Crypto inverse-power options
In this section, we present characteristic function-based pricing–hedging formulas for inverse and Quanto inverse contracts on cryptos and their power-type extensions. [^12] The idea of treating inverse options as special cases of negative-powered options is in fact motivated from an independent study concerning Poisson functionals ([Xia, 2022] [59]). For example, a standard inverse option (pre-conversion) can be regarded as a power option on the same underlying with (negative) power $-1$. As aforementioned, for USD-denominated investors, an inverse contract is substantially equivalent to a direct contract upon crypto/USD conversion at expiry, whose valuation will follow the same procedures for standard derivatives. A Quanto inverse contract, on the other hand, serves to mitigate currency risk borne by the inverse contract via specifying a fixed conversion rate at inception. Thus, the contract remains “inverse” in nature and will exhibit a (locally) concave payoff for (long) call options as shown in [Alexander et al., 2023, Fig. 4] [2].
@@ 3.1 Inverse-power options
We begin by looking at the terminal payoffs of inverse-power options on $S$ expressed in units of cryptos. The terminal payoff of an inverse-power call option is given by [^13] In what follows the subscripts “(ip),” “(p),” “(qi),” and “(qip)” stand for “inverse-power,” “power,” “Quanto inverse,” and “Quanto inverse-power,” respectively.
@@
with original strike price $K$ (\$), expiry date $T$, and two predetermined power coefficients $p_{1},p_{2}\geq 0$; for a similar put option, the payoff is
@@
This power mechanism provides the option investor with flexible nonlinear leverage on the crypto/USD conversion rate in the same way as usual power options, and from the above transformations, the inverse-power call option is equivalent to a put option on a powered USD/crypto conversion rate, and similarly for the put option. The power coefficient $p_{2}$ is explicitly used to ensure that the powered prices are at the same scale and hence ease comparison, though each $K^{p_{2}}$ can be treated as a new strike price. The actual impact of $p_{1}$ on risk exposure depends on the magnitude of the crypto/USD conversion rate relative to 1 as a threshold. For instance, as per status quo, if $S$ is for Bitcoin or Ethereum, then $p_{1}>1$ is most likely associated with increased levels of exposure while $p_{1}<1$ imposes a decreasing effect; however, if $S$ is for Ripple or Dogecoin (with options also being actively traded), this effect should be completely reversed. Therefore, the choice of the powers in designing these contracts should be made on a case-by-case basis, and for the same reason, the power mechanism will be ineffective for cryptos pegged to USD at 1-to-1, such as Tether.
@@
As discussed in [Alexander et al., 2023] [2], for USD-denominated investors, the above payoffs can be translated into USD at expiry, making the true payoff function approximately equal to that of the direct counterparts. In the call option case, this means that
@@
In contrast, the currency conversion can be realized through a predetermined rate $R$ (\$), which then gives rise to the following Quanto inverse-type payoff:
@@
which, mathematically, can be viewed as a scalar multiple of the inverse-power option payoff without currency conversion. This in turn shows that the pricing problem of Quanto inverse-power options from a USD-denominated investor’s perspective is analogous to that of inverse-power options for a non-USD-based investor.
For the empirical analysis in Section 5 which exclusively involves call options, we will use the true (approximate) payoff to USD-denominated investors in (3.1.3). The value of this direct call option at time $t_{0}\in[0,T)$ is given by the formula due to [Bakshi and Madan, 2000] [5]
@@
$\varPi_{1}=\frac{1}{2}+\frac{1}{\pi}\int^{\infty}_{0}\mathrm{Re}\bigg{[}\frac{K^{-{\rm i}u}\phi_{\log S_{T}|t_{0}}(u-{\rm i})}{{\rm i}u\phi_{\log S_{T}|t_{0}}(-{\rm i})}\bigg{]}{\rm d}u,\quad\varPi_{2}=\frac{1}{\pi}+\frac{1}{\pi}\int^{\infty}_{0}\mathrm{Re}\bigg{[}\frac{K^{-{\rm i}u}\phi_{\log S_{T}|t_{0}}(u)}{{\rm i}u}\bigg{]}{\rm d}u.$ (3.1.6)
@@
In spite of these well-known formulas, for our calibration purposes we choose to work with the parity-implied formula
@@
For the Quanto inverse-power options (again with USD-denomination), we present general characteristic function-based pricing formulas in the next proposition, which accommodate a broad range of underlying models. Similar to their direct option counterparts (see, e.g., [Xia, 2017, Sect. 5] [56]), these formulas are very amenable to efficient implementations and elude the need for Monte Carlo simulations.
@@
In the above formula, the quantity $\phi_{\log S_{T}|t_{0}}({\rm i}p_{1})=\mathsf{E}_{t_{0}}\big{[}S^{-p_{1}}_{T}\big{]}$, assumed finite, represents the contemporaneous value of an “inverse-power forward” on $S$, or a futures contract on the USD/crypto rate. As the term suggests, such a forward is not to be confused with an inverse futures, the latter being the inverse of crypto/USD futures rates.
The pricing formulas for usual Quanto inverse options are immediately obtained by setting $p_{1}=p_{2}=1$ in (3.1.9) and (3.1.10), and we have (in USD)
@@ 3.2 Dynamic hedging
It is widely known that perfect replication of options is possible when the underlying price is driven by diffusions or Poisson-type jumps. Direct implications on market incompleteness notwithstanding, in the presence of fractional stochastic volatility and price–volatility co-jumps, we aim to construct partial hedges for the Quanto inverse-power options following Proposition 2, by utilizing the underlying crypto/USD conversion rate ($S$) and the associated variance swap ($V_{S}$), both being investible at present.
@@
Here, recall that $V_{S}(t,T)$, for every $t$, represents the initial value of a (different) variance swap expiring at time $T$, rather than the time-$t$ value of the same variance swap issued at time 0. Hence, the necessary hedging instruments for the Quanto inverse-power option consist of the spot conversion rate and a flow of variance swaps available prior to $T$, the latter being equivalent to the collection of forward variance curves, $\{(\partial_{T}V_{S}(v,T))_{v\in[t,T)}:t\in[0,T)\}$, which is well-defined (see the proof of Corollary 1 in Appendix A), parallel to the case of the rough Heston model ([El Euch and Rosenbaum, 2018] [20]).
@@
As a few remarks, on the right side of (1), the jump components $\varDelta S_{s}$, $\varDelta V_{S}(s,T)$, and $\varDelta C^{\rm(qip)}_{s}$ all contain non-hedge-able jump risks present in both crypto prices and volatility. Also, since the precise activity rate $A$ is not observed, [^14] For the same reason, integrated variance is not directly observable, despite being estimable from corresponding time series. unless $X$ is purely discontinuous ($\sigma_{X}=0$), the quadratic variation term $\langle S^{\rm c},S^{\rm c}\rangle_{s}$ will induce additional volatility risk that cannot be perfectly hedged. Still, this term may be estimated by resorting to compound Poisson approximation techniques which truncate the jump space and coalesce small jumps into the Brownian component; we refer to [Schoutens, 2003, Sect. 8.2.1] [47] for details. [^15] However, note that such estimation should be done under some (induced) physical measure $\mathsf{P}$ and should go through the characteristic changes specified in (2.2). All the other terms in (1) constitute perfect hedges given the accessibility to $S$ and $V_{S}$. In particular, while the integral with respect to $S$ is simply a delta hedge requiring periodic investments in the underlying crypto (with USD), the integral with respect to $V_{S}$, being a Vega hedge, consists in buying and selling (or shorting and buying back) newly issued variance swaps over consecutive periods, whose practical validity should depend on the frequency of issuance.
Moreover, an alternative hedge, which is less complete but easier to implement, is to trade a single variance swap $U$ issued at time 0 and expiring at $T$ instead of the flow of variance swaps mentioned above. Put precisely, since $C^{\rm(qip)}$ is by construction a $\mathsf{Q}$-martingale and so is the value process of this single swap, namely
@@ 4 Specification analysis
We now attend to the issue of specifying the fractional stochastic-volatility model proposed in Section 2 for efficient implementation. We will consider special forms of the fractional kernel $h$ in (2.1.3) according to the discussions in [Wang and Xia, 2022, Sect. 2.2] [53], all of which have the same parametrization $\{\kappa>0,d\in(1/2,1]\}$. We will also adopt some of the new Lévy models developed in [Fei and Xia, 2024] [22] for the background-driving processes $X$ and $Y$. Again, the main goal is to select a suitable class of models that capture the key traits of crypto price dynamics, including inherent price jumps, volume impact, stochastic volatility with short-range dependence, as well as volatility jumps of suitable activity and their co-movements with the prices. The analysis will focus on obtaining various explicit expressions for the ingredients present in the pricing–hedging formulas.
@@ 4.1 Fractional kernels and activity rate
$J_{1}(t-t_{0}):=\int^{t}_{t_{0}}H_{1}(t-s){\rm d}s=\frac{\kappa\mathrm{\Gamma}(d)(t-t_{0})-(\mathrm{\Gamma}(d+1)+\kappa(t-t_{0})\mathrm{\Gamma}(d,\kappa(t-t_{0}))-\mathrm{\Gamma}(d+1,\kappa(t-t_{0})))}{\kappa^{d+1}\mathrm{\Gamma}(d)},$
@@
The type-II kernel that appeared in [Wang and Xia, 2022] [53] was designed to match the usual Ornstein–Uhlenbeck process driven by a fractional process with the Riemann–Liouville kernel. In particular, by choosing
@@
The two pieces subtly separate the rough behavior of the activity rate at small time intervals from its mean reversion in the long run, with the long-run average being $\lim_{t\rightarrow\infty}\mathsf{E}[A_{t}]=m+\mathsf{E}[Y_{1}]((1-d)/\kappa)^{d}/((1-d)\mathrm{\Gamma}(d+1))$. The tail-integrated kernel is given by
@@ 4.2 Regulating kernels and base processes
$\bigg{(}\mu_{X}=\frac{\lambda_{X}(1-\eta^{2})}{b_{X}\eta},\sigma_{X},\nu_{X}({\rm d}z)=\lambda_{X}\bigg{(}\frac{b_{X}\eta}{\eta^{2}+1}\bigg{)}e^{-b_{X}(z\;\mathrm{sgn}z)\eta^{\mathrm{sgn}z}}{\rm d}z\bigg{)},\quad(0,0,\nu_{Y}({\rm d}z)=\lambda_{Y}b_{Y}e^{-b_{Y}z}\mathds{1}_{\{z>0\}}{\rm d}z).$
@@
$\phi_{Z^{(n)}_{1}}(u)=\begin{cases}\displaystyle\exp\bigg{(}a_{X}b^{c_{X}}_{X}\mathrm{\Gamma}(-c_{X})\bigg{(}{}_{2}\mathrm{F}_{1}\bigg{(}-c_{X},\frac{1}{n};\frac{1}{n}+1;\frac{{\rm i}u}{b_{X}\mathrm{\Gamma}(n+1)}\bigg{)}-1\bigg{)}\bigg{)}\;&\text{if }c_{X}>0\\ \displaystyle\bigg{(}1-\frac{{\rm i}u}{b_{X}\mathrm{\Gamma}(n+1)}\bigg{)}^{-a_{X}}\exp\bigg{(}-\frac{{\rm i}a_{X}nu}{b_{X}\mathrm{\Gamma}(n+2)}{}_{2}\mathrm{F}_{1}\bigg{(}1,\frac{1}{n}+1;\frac{1}{n}+2;\frac{{\rm i}u}{b_{X}\mathrm{\Gamma}(n+1)}\bigg{)}\bigg{)}\;&\text{if }c_{X}=0,\end{cases}$
@@
$\qquad\nu_{X}({\rm d}z)=\int^{t}_{0}{\rm d}w\;\frac{a_{X}b^{c_{X}}_{X}(b_{X}\mathrm{\Gamma}(n+1)w)^{1/n}\mathrm{\Gamma}(-c_{X}-1/n,b_{X}\mathrm{\Gamma}(n+1)w)\gamma_{\theta w,w}({\rm d}z)}{nw}\bigg{)},$
@@
where $\mu_{X}=\mathsf{E}[X_{1}]$ is computed in the proof of Proposition 3 in Appendix A and $\gamma_{\theta w,w}$ denotes the Gaussian measure with mean $\theta w$ and variance $w$.
All of the above specified characteristic functions for $X$ and $Y$ are to be put into the conditional characteristic function in (1), after which a pricing engine can be initiated with up to two (parallel) numerical integrals to be computed. With any of the three types of fractional kernels in Section 4.1, the integral of $H$ has been explicitly given, which leaves us with only one numerical integral. This undoubtedly lays the foundation for efficient pricing–hedging. To go one step further, the next proposition, serving as the main result of this section, gives a perfectly analytical formula for the conditional characteristic function in when $h$ happens to be the type-III (fractional) kernel. For a succinct presentation, let us introduce the following notations (for generic $u$):

### v3

@@ 3 Crypto inverse-power options
As noted in [Alexander et al., 2023] [2], approximately 90% of active crypto options trading currently takes place on the Deribit exchange. A distinctive feature of Deribit’s options is their inverse nature: Contracts are denominated and settled in the underlying crypto although the underlying asset is quoted in USD. Deribit reports USD-equivalent prices by multiplying the crypto-settled payoff by the crypto/USD conversion rate. For example, a Bitcoin option’s profit or loss is calculated and settled in Bitcoin, even though the option prices and strike prices are displayed in USD. [^12] For a comprehensive explanation of Deribit’s inverse contract structure and profit calculation, refer to their official guide here.
For USD-denominated investors, such an inverse contract is substantially equivalent to a direct contract upon immediate crypto/USD conversion at expiry, whose valuation will follow the same procedures for standard derivatives. In this regard, the terminal payoff of an inverse call option written on $S$ with strike price $K$ (\$) and expiry date $T$ satisfies that [^13] In what follows the subscripts “(i),” “(ip),” “(p),” “(qi),” and “(qip)” stand for “inverse,” “inverse-power,” “power,” “Quanto inverse,” and “Quanto inverse-power,” respectively.
[…]
where, upon currency conversion, $C_{T}$ is the payoff of a similar (direct) standard call option. A Quanto inverse contract, on the other hand, serves to mitigate currency risk borne by the inverse contract via specifying a fixed conversion rate $R$ (crypto/USD) at inception; its terminal payoff is given by
$C^{\rm(qi)}_{T}=RC^{\rm(i)}_{T}=R\bigg{(}1-\frac{K}{S_{T}}\bigg{)}^{+}\;(\$).$ (3.2)
Thus, the contract remains “inverse” in nature and can exhibit a locally concave payoff for (long) calls as shown in [Alexander et al., 2023, Fig. 4] [2].
In the rest of this section, we present characteristic function-based pricing–hedging formulas for these Quanto inverse contracts as well as their power-type extensions, [^14] The idea of treating inverse options as special cases of negative-powered options is motivated from an independent study concerning Poisson functionals ([Xia, 2022] [58]). As seen from (3.1), a standard inverse option (pre-conversion) can be regarded as a power option on the same underlying with (negative) power $-1$. which provide significant flexibility for adjusting risk exposure.
@@ 3.1 Inverse-power options
The terminal payoff of an inverse-power call option on $S$, in the setting of (3.1), is given by
@@
with two predetermined power coefficients $p_{1},p_{2}\geq 0$; for a similar put option, the payoff is
@@
This power mechanism provides the option investor with nonlinear leverage on the crypto/USD conversion rate in the same way as usual power options. From the above transformations, the inverse-power call option is equivalent to a put option on a powered USD/crypto conversion rate, and similarly for the put option. The power coefficient $p_{2}$ is explicitly used to ensure that the powered prices are at the same scale and hence ease comparison, though each $K^{p_{2}}$ can be treated as a new strike price. The actual impact of $p_{1}$ on risk exposure depends on the magnitude of the crypto/USD conversion rate relative to 1 as a threshold. For instance, as per status quo, if $S$ is for Bitcoin or Ethereum, then $p_{1}>1$ is most likely associated with increased levels of exposure while $p_{1}<1$ imposes a decreasing effect; however, if $S$ is for Dogecoin or Polygon (MATIC) (also with options available), this effect should be completely reversed. Therefore, the choice of the powers in designing these contracts should be made on a case-by-case basis; for the same reason, the power mechanism will be ineffective for cryptos pegged to USD at 1-to-1, such as Tether.
@@
As discussed earlier, for USD-denominated investors, the above payoffs can be translated into USD at expiry, making the true payoff function (approximately) equal to that of the direct counterparts. In a similar manner to (3.1), this means that
@@
In contrast, as done in (3.2), leveraging a predetermined conversion rate $R$ gives rise to the following Quanto inverse-type payoff:
@@
which is technically a scalar multiple of the inverse-power option payoff without conversion. This in turn shows that the pricing problem of Quanto inverse-power options from a USD-denominated investor’s perspective is analogous to that of inverse-power options for a non-USD-based investor.
For the empirical analysis in Section 5 exclusively involving call options, we will use the (approximate) payoff to USD-denominated investors in (3.1.3), in line with Deribit’s option quotes. The value of this direct call option at time $t_{0}\in[0,T)$ is given by the formula due to [Bakshi and Madan, 2000] [5]
@@
$\varPi_{1}=\frac{1}{2}+\frac{1}{\pi}\int^{\infty}_{0}\mathrm{Re}\bigg{[}\frac{K^{-{\rm i}u}\phi_{\log S_{T}|t_{0}}(u-{\rm i})}{{\rm i}u\phi_{\log S_{T}|t_{0}}(-{\rm i})}\bigg{]}{\rm d}u,\quad\varPi_{2}=\frac{1}{2}+\frac{1}{\pi}\int^{\infty}_{0}\mathrm{Re}\bigg{[}\frac{K^{-{\rm i}u}\phi_{\log S_{T}|t_{0}}(u)}{{\rm i}u}\bigg{]}{\rm d}u.$ (3.1.6)
@@
Despite these well-known formulas, for our calibration purposes we choose to work with the parity-implied formula
@@
For the Quanto inverse-power options (again, with USD-denomination), we give general characteristic function-based pricing formulas in the next proposition, accommodating a broad range of underlying models. Similar to their direct option counterparts (see, e.g., [Xia, 2017, Sect. 5] [55]), these formulas are very amenable to efficient implementations and elude the need for Monte Carlo simulations.
@@
In the above formula, the quantity $\phi_{\log S_{T}|t_{0}}({\rm i}p_{1})=\mathsf{E}_{t_{0}}\big{[}S^{-p_{1}}_{T}\big{]}$, assumed finite, represents the contemporaneous value of an “inverse-power forward” on $S$, or a futures contract on the powered USD/crypto rate. As the term suggests, such a forward is not to be confused with an inverse futures, the latter being the inverse of crypto/USD futures rates.
The pricing formulas for usual Quanto inverse options (recalling (3.2) e.g.) are immediately obtained by setting $p_{1}=p_{2}=1$ in (3.1.9) and (3.1.10), and we have (in USD)
@@ 3.2 Dynamic hedging
It is widely known that perfect replication of options is possible when the underlying price is driven by diffusions or Poisson-type jumps. Direct implications on market incompleteness notwithstanding, given stochastic volatility and price–volatility co-jumps, we aim to construct partial hedges for the Quanto inverse-power options following Proposition 2, by utilizing the underlying crypto/USD conversion rate ($S$) and the associated variance swap ($V_{S}$), both being investible at present.
@@
Here, recall that $V_{S}(t,T)$, for every $t$, represents the initial value of a (different) variance swap expiring at time $T$, rather than the time-$t$ value of the same variance swap issued at time 0. Hence, the necessary hedging instruments for the Quanto inverse-power option consist of the spot conversion rate and a flow of variance swaps available prior to $T$, the latter being equivalent to the collection of forward variance curves, $\{(\partial_{T}V_{S}(v,T))_{v\in[t,T)}:t\in[0,T)\}$, which is well-defined (see the proof of Corollary 1 in A), parallel to the case of the rough Heston model ([El Euch and Rosenbaum, 2018] [20]).
@@
On the right side of (1), the jump components $\varDelta S_{s}$, $\varDelta V_{S}(s,T)$, and $\varDelta C^{\rm(qip)}_{s}$ all contain non-hedge-able jump risks present in both crypto prices and volatility. Also, since the precise activity rate $A$ is not observed, [^15] For the same reason, integrated variance is not directly observable, despite being estimable from corresponding time series. unless $X$ is purely discontinuous ($\sigma_{X}=0$), the quadratic variation term $\langle S^{\rm c},S^{\rm c}\rangle_{s}$ will induce additional volatility risk that cannot be perfectly hedged. Still, this term may be estimated by resorting to compound Poisson approximation techniques which truncate the jump space and coalesce small jumps into the Brownian component; we refer to [Schoutens, 2003, Sect. 8.2.1] [46] for details. [^16] However, note that such estimation should be done under some (induced) physical measure $\mathsf{P}$ and should go through the characteristic changes specified in (2.2). All the other terms in (1) constitute perfect hedges given the accessibility to $S$ and $V_{S}$. In particular, while the integral with respect to $S$ is simply a delta hedge requiring periodic investments in the underlying crypto (with USD), the integral with respect to $V_{S}$, being a vega hedge, consists in buying and selling (or shorting and buying back) newly issued variance swaps over consecutive periods, whose practical validity should depend on the frequency of issuance.
An alternative hedge, which is less complete but easier to implement, is to trade a single variance swap $U$ issued at time 0 and expiring at $T$ instead of the flow of variance swaps mentioned above. Put precisely, since $C^{\rm(qip)}$ is by construction a $\mathsf{Q}$-martingale and so is the value process of this single swap, namely
@@ 4 Specification analysis
We now attend to the issue of specifying the fractional stochastic-volatility model proposed in Section 2 for efficient implementation. We will consider special forms of the fractional kernel $h$ in (2.1.3) according to the discussions in [Wang and Xia, 2022, Sect. 2.2] [52], all of which have the same parametrization $\{\kappa>0,d\in(1/2,1]\}$. We will also adopt some of the new Lévy models developed in [Fei and Xia, 2024] [22] for the background-driving processes $X$ and $Y$. Again, the main goal is to select a suitable class of models that capture the key traits of crypto price dynamics, including inherent price jumps, volume impact, stochastic volatility with short-range dependence, as well as volatility jumps of suitable activity and their co-movements with the prices. The analysis will focus on obtaining various explicit expressions for the ingredients in the pricing–hedging formulas.
@@ 4.1 Fractional kernels and activity rate
$J_{1}(t-t_{0})$ $:=\int^{t}_{t_{0}}H_{1}(t-s){\rm d}s$
$=\frac{\kappa\mathrm{\Gamma}(d)(t-t_{0})-(\mathrm{\Gamma}(d+1)+\kappa(t-t_{0})\mathrm{\Gamma}(d,\kappa(t-t_{0}))-\mathrm{\Gamma}(d+1,\kappa(t-t_{0})))}{\kappa^{d+1}\mathrm{\Gamma}(d)},$
@@
The type-II kernel appearing in [Wang and Xia, 2022] [52] was designed to match the usual Ornstein–Uhlenbeck process driven by a fractional process with the Riemann–Liouville kernel. In particular, by choosing
@@
The two pieces subtly separate the rough behavior of the activity rate at small time intervals from its mean-reverting behavior in the long run, with the long-run average being $\lim_{t\rightarrow\infty}\mathsf{E}[A_{t}]=m+\mathsf{E}[Y_{1}]((1-d)/\kappa)^{d}/((1-d)\mathrm{\Gamma}(d+1))$. The tail-integrated kernel is given by
@@ 4.2 Regulating kernels and base processes
$\bigg{(}\mu_{X}=\frac{\lambda_{X}(1-\eta^{2})}{b_{X}\eta},\sigma_{X},\nu_{X}({\rm d}z)=\lambda_{X}\bigg{(}\frac{b_{X}\eta}{\eta^{2}+1}\bigg{)}e^{-b_{X}(z\;\mathrm{sgn}z)\eta^{\mathrm{sgn}z}}{\rm d}z\bigg{)},$
$(0,0,\nu_{Y}({\rm d}z)=\lambda_{Y}b_{Y}e^{-b_{Y}z}\mathds{1}_{\{z>0\}}{\rm d}z).$
@@
$\phi_{Z^{(n)}_{1}}(u)=\begin{cases}\displaystyle\exp\bigg{(}a_{X}b^{c_{X}}_{X}\mathrm{\Gamma}(-c_{X})\bigg{(}{}_{2}\mathrm{F}_{1}\bigg{(}-c_{X},\frac{1}{n};\frac{1}{n}+1;\frac{{\rm i}u}{b_{X}\mathrm{\Gamma}(n+1)}\bigg{)}-1\bigg{)}\bigg{)}\\ \qquad\text{if }c_{X}>0\\ \displaystyle\bigg{(}1-\frac{{\rm i}u}{b_{X}\mathrm{\Gamma}(n+1)}\bigg{)}^{-a_{X}}\exp\bigg{(}-\frac{{\rm i}a_{X}nu}{b_{X}\mathrm{\Gamma}(n+2)}{}_{2}\mathrm{F}_{1}\bigg{(}1,\frac{1}{n}+1;\frac{1}{n}+2;\frac{{\rm i}u}{b_{X}\mathrm{\Gamma}(n+1)}\bigg{)}\bigg{)}\\ \qquad\text{if }c_{X}=0,\end{cases}$
@@
$\qquad\nu_{X}({\rm d}z)=\int^{\infty}_{0}{\rm d}w\;\frac{a_{X}b^{c_{X}}_{X}(b_{X}\mathrm{\Gamma}(n+1)w)^{1/n}\mathrm{\Gamma}(-c_{X}-1/n,b_{X}\mathrm{\Gamma}(n+1)w)\gamma_{\theta w,w}({\rm d}z)}{nw}\bigg{)},$
@@
where $\mu_{X}=\mathsf{E}[X_{1}]$ is computed in the proof of Proposition 3 in A and $\gamma_{\theta w,w}$ denotes the Gaussian measure with mean $\theta w$ and variance $w$.
All of the above specified characteristic functions for $X$ and $Y$ are to be put into the conditional characteristic function in (1), after which a pricing engine can be initiated with up to two (parallel) numerical integrals to be computed. With any of the three types of fractional kernels in Section 4.1, the integral of $H$ has been explicitly given, which leaves us with only one numerical integral. This undoubtedly lays the foundation for efficient pricing–hedging. To go one step further, the next proposition, serving as the main result of this section, gives a perfectly analytical formula for the conditional characteristic function when $h$ happens to be the type-III (fractional) kernel. For a succinct presentation, let us introduce the following notations (for generic $u$):